# FlashAttention-3: Fast and Accurate Attention with Asynchrony and Low-precision

### v1

@@ 1 Introduction
In this work, we build on the work of Dao et al. [17] on developing exact-attention algorithms that integrate knowledge of the GPU’s execution model and hardware characteristics into their high-level design. In [17], Dao et al. introduced FlashAttention, a novel tiling strategy for parallelizing attention that eliminates intermediate reads/writes to slow global memory through fusing all of the attention operations into a single GPU kernel. Dao [15] restructured the algorithm as FlashAttention-2 to also parallelize over the sequence length dimension and perform the inner loop of the forward pass over blocks of the key and value matrices, thus improving the occupancy and distribution of work on the GPU. However, we observe that FlashAttention-2 nonetheless achieves poor utilization on newer GPUs relative to optimized matrix-multiplication (GEMM) kernels, such as 35% vs. 80-90% on the Hopper H100 GPU. Partially, this may be attributed to implementation-level differences, such as not using Hopper-specific instructions in place of Ampere ones when targeting the Tensor Cores.
@@ 2.2 GPU hardware characteristics and execution model
GPUs are throughput processors that rely on concurrency and asynchrony to hide memory and execution latencies. For async memory copy between GMEM and SMEM, Hopper has the Tensor Memory Accelerator (TMA) as a dedicated hardware unit [38, §7.29]. Furthermore, unlike prior architectures such as Ampere, the Tensor Core of Hopper, exposed via the warpgroup-wide WGMMA instruction [39, §9.7.14], is also asynchronous and can source its inputs directly from shared memory.
Hardware support for asynchrony allows for warp-specialized kernels, where the warps of a CTA are divided into producer or consumer roles that only ever issue either data movement or computation. Generically, this improves the compiler’s ability to generate optimal instruction schedules [4]. In addition, Hopper supports the dynamic reallocation of registers between warpgroups via setmaxnreg [39, §9.7.17.1], so those warps doing MMAs can obtain a larger share of RMEM than those just issuing TMA (for which only a single thread is needed).
@@ 3.3 Low-precision with FP8
Instead, for FP8 FlashAttention-3 we opt for option (2). For the in-kernel transpose, we take advantage of the LDSM (ldmatrix) and STSM (stmatrix) instructions, which involve a warp of threads collectively loading SMEM to RMEM and storing RMEM to SMEM at a granularity of 128 bytes. [^8] In the PTX documentation, LDSM/STSM are described as copying $8\times 8$ matrices with 16-bit entries [39, §9.7.13.4.15-16], but we can pack 8-bit entries two at a time to use LDSM/STSM in the context of FP8 precision. However, the transpose versions of LDSM/STSM cannot split packed 8-bit entries, which necessitates certain register movements in between LDSM and STSM to actually perform a tile-wise transpose; we omit the details. The LDSM/STSM instructions are both register efficient, allowing us to execute them in the producer warpgroup, and capable of transposing layouts when doing memory copy. Moreover, after the first iteration we can arrange for the transpose of the next $\mathbf{V}$ tile to be executed in the shadow of the two WGMMAs that involve the preceding $\mathbf{V}$ and current $\mathbf{K}$ tile.
@@
    Incoherent processing: to even out outliers, we multiply $\mathbf{Q}$ and $\mathbf{K}$ with a random orthogonal matrix $\mathbf{M}$ before quantizing to FP8. Since $\mathbf{M}$ is orthogonal, $\mathbf{M}\mathbf{M}^{\top}=I$ and so $(\mathbf{Q}\mathbf{M})(\mathbf{K}\mathbf{M})^{\top}=\mathbf{Q}\mathbf{K}^{\top}$, i.e., multiplying both $\mathbf{Q}$ and $\mathbf{K}$ with $\mathbf{M}$ does not change the attention output. This serves to “spread out” the outliers since each entry of $\mathbf{Q}\mathbf{M}$ or $\mathbf{K}\mathbf{M}$ is a random sum of entries of $\mathbf{Q}$ or $\mathbf{K}$, thus reducing quantization error. In practice, we follow Chee et al. [9] and choose $\mathbf{M}$ to be the product of random diagonal matrices of ±$\pm 1$ and a Hadamard matrix, which can be multiplied in $O(d\log d)$ instead of $O(d^{2})$, and can also be fused with the rotary embedding at no extra computation cost.
@@ Acknowledgments
We are grateful to the NVIDIA CUTLASS team (especially Haicheng Wu, Aniket Shivam, and Cris Cecka) for helping us understand Hopper’s programming model and for their library, which provides clean and powerful building blocks for the implementation of FlashAttention-3. We thank the cuDNN team for the idea of in-kernel transpose for FP8. We appreciate Driss Guessous for integrating FlashAttention to PyTorch. FlashAttention-3 has benefited from insightful discussions with Horace He on different attention variants, with Hao Liu and Phil Wang on distributed attention, and with Daniel Haziza and Chris De Sa on quantization. We thank Meta, Together AI, and Princeton Language and Intelligence (PLI) for compute support.
@@ Appendix A Related Work
Quantization is a promising approach to speed up attention, but they have mostly focused on reducing the space for KV cache for inference efficiency. QuIP [9] uses incoherent processing to reduce the quantization, and we adapted this technique for FP8 FlashAttention-3. Recent work suggests that for inference the KV cache is highly compressible down to 4-, 3-, or even 2-bits [26, 33]. However, quantization during training is still challenging as higher precision is typically required for stable training.

### v2

@@ 1 Introduction
In this work, we build on the work of Dao et al. [17] on developing exact-attention algorithms that integrate knowledge of the GPU’s execution model and hardware characteristics into their high-level design. In [17], Dao et al. introduced FlashAttention, a novel tiling strategy for parallelizing attention that eliminates intermediate reads/writes to slow global memory through fusing all of the attention operations into a single GPU kernel. Dao [15] restructured the algorithm as FlashAttention-2 to also parallelize over the sequence length dimension and perform the inner loop of the forward pass over blocks of the key and value matrices, thus improving the occupancy and distribution of work on the GPU. However, we observe that FlashAttention-2 nonetheless achieves poor utilization on newer GPUs relative to optimized matrix-multiplication (GEMM) kernels, such as 35% vs. 80-90% on the Hopper H100 GPU. Partially, this may be attributed to implementation-level differences, such as not using Hopper-specific instructions in place of Ampere ones when targeting the Tensor Cores. Several work such as ThunkerKitten [52] and cuDNN 9 [39] has shown that with Hopper-specific instructions and tile-based abstractions, one can speedup attention computation and simplify the implementation.
@@ 2.2 GPU hardware characteristics and execution model
GPUs are throughput processors that rely on concurrency and asynchrony to hide memory and execution latencies. For async memory copy between GMEM and SMEM, Hopper has the Tensor Memory Accelerator (TMA) as a dedicated hardware unit [38, §7.29]. Furthermore, unlike prior architectures such as Ampere, the Tensor Core of Hopper, exposed via the warpgroup-wide WGMMA instruction [40, §9.7.14], is also asynchronous and can source its inputs directly from shared memory.
Hardware support for asynchrony allows for warp-specialized kernels, where the warps of a CTA are divided into producer or consumer roles that only ever issue either data movement or computation. Generically, this improves the compiler’s ability to generate optimal instruction schedules [4]. In addition, Hopper supports the dynamic reallocation of registers between warpgroups via setmaxnreg [40, §9.7.17.1], so those warps doing MMAs can obtain a larger share of RMEM than those just issuing TMA (for which only a single thread is needed).
@@ 3.3 Low-precision with FP8
Instead, for FP8 FlashAttention-3 we opt for option (2). For the in-kernel transpose, we take advantage of the LDSM (ldmatrix) and STSM (stmatrix) instructions, which involve a warp of threads collectively loading SMEM to RMEM and storing RMEM to SMEM at a granularity of 128 bytes. [^8] In the PTX documentation, LDSM/STSM are described as copying $8\times 8$ matrices with 16-bit entries [40, §9.7.13.4.15-16], but we can pack 8-bit entries two at a time to use LDSM/STSM in the context of FP8 precision. However, the transpose versions of LDSM/STSM cannot split packed 8-bit entries, which necessitates certain register movements in between LDSM and STSM to actually perform a tile-wise transpose; we omit the details. The LDSM/STSM instructions are both register efficient, allowing us to execute them in the producer warpgroup, and capable of transposing layouts when doing memory copy. Moreover, after the first iteration we can arrange for the transpose of the next $\mathbf{V}$ tile to be executed in the shadow of the two WGMMAs that involve the preceding $\mathbf{V}$ and current $\mathbf{K}$ tile.
@@
    Incoherent processing: to even out outliers, we multiply $\mathbf{Q}$ and $\mathbf{K}$ with a random orthogonal matrix $\mathbf{M}$ before quantizing to FP8. Since $\mathbf{M}$ is orthogonal, $\mathbf{M}\mathbf{M}^{\top}=I$ and so $(\mathbf{Q}\mathbf{M})(\mathbf{K}\mathbf{M})^{\top}=\mathbf{Q}\mathbf{K}^{\top}$, i.e., multiplying both $\mathbf{Q}$ and $\mathbf{K}$ with $\mathbf{M}$ does not change the attention output. This serves to “spread out” the outliers since each entry of $\mathbf{Q}\mathbf{M}$ or $\mathbf{K}\mathbf{M}$ is a random sum of entries of $\mathbf{Q}$ or $\mathbf{K}$, thus reducing quantization error. In practice, we follow Chee et al. [9] and Tseng et al. [58] and choose $\mathbf{M}$ to be the product of random diagonal matrices of ±$\pm 1$ and a Hadamard matrix, which can be multiplied in $O(d\log d)$ instead of $O(d^{2})$, and can also be fused with the rotary embedding at no extra computation cost.
@@ Acknowledgments
We are grateful to the NVIDIA CUTLASS team (especially Haicheng Wu, Aniket Shivam, and Cris Cecka) for helping us understand Hopper’s programming model and for their library, which provides clean and powerful building blocks for the implementation of FlashAttention-3. We thank the cuDNN team for the idea of in-kernel transpose for FP8. The idea of overlapping GEMMs and softmax was inspired by insightful conversations with Christopher Ré, Benjamin Spector, Aniket Shivam, and Markus Hoehnerbach. The pingpong scheduling is adapted from the warp-specialized pingpong GEMM implementation in CUTLASS. We appreciate Driss Guessous for integrating FlashAttention to PyTorch. FlashAttention-3 has benefited from helpful discussions with Horace He on different attention variants, with Hao Liu and Phil Wang on distributed attention, and with Daniel Haziza and Chris De Sa on quantization. We thank Meta, Together AI, and Princeton Language and Intelligence (PLI) for compute support.
@@ Appendix A Related Work
Quantization is a promising approach to speed up attention, but they have mostly focused on reducing the space for KV cache for inference efficiency. QuIP [9] and QuIP#[58] use incoherent processing to reduce the quantization, and we adapted this technique for FP8 FlashAttention-3. Recent work suggests that for inference the KV cache is highly compressible down to 4-, 3-, or even 2-bits [26, 33]. However, quantization during training is still challenging as higher precision is typically required for stable training.